# A refined Weyl character formula for comodules on $\operatorname{GL}(2,A)$

### v2

@@ 2. The adjoint representation of $\operatorname{GL}(2,A)$ and $\operatorname{SL}(2,A)$
The adjoint representation for $\operatorname{Lie}(\operatorname{SL}(2,A))$.
Let $\mathfrak{sl}(2,A):=A\{e_{12},e_{11}-e_{22},e_{21}\}\subseteq\mathfrak{gl}(2,A)$, and let $x:=e_{12},H:=e_{11}-e_{22},y:=e_{21}$ and let $B(s):=\{x,H,y\}$. Define for any commutative $A$-algebra $C$ the following: Let $B(s)_{C}:=\{x\otimes 1_{C},H\otimes 1_{C},y\otimes 1_{C}\}$ be a basis for $\mathfrak{sl}(2,A)\otimes_{A}C$. Here $1_{C}\in C$ is the multiplicative unit. You may check the following:
@@
Hence $\mathfrak{sl}(2,A)\subseteq\mathfrak{gl}(2,A)$ is a sub $k[G]$-comodule. Let $S:=\operatorname{SL}(2,A):=V(D-1)\subseteq\operatorname{GL}(2,A)$. we get the induced $k[S]$-comodule
$\operatorname{Ad}(S):\mathfrak{sl}(2,A)\rightarrow\mathfrak{sl}(2,A)\otimes_{A}k[S]$
@@
There is an isomorphism $\operatorname{Lie}(\operatorname{SL}(2,A))\cong\mathfrak{sl}(2,A)$. Let $\operatorname{Lie}(T):=A\{H\}$ and let $\operatorname{Lie}(\operatorname{SL}(2,A))_{R_{S}}:=A\{x,y\}$. It follow there are comodule structures
@@
$\Delta_{2}:\operatorname{Lie}(\operatorname{SL}(2,A))_{R_{S}}\rightarrow\operatorname{Lie}(\operatorname{SL}(2,A))_{R_{S}}\otimes k[N]$
@@
$\operatorname{Lie}(\operatorname{SL}(2,A))\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{SL}(2,A))_{R_{S}}.$
@@
The set $R_{S}:=\{-2,2\}$ is the roots of $\operatorname{SL}(2,A)$ and the direct sum decomposition
$\operatorname{Lie}(\operatorname{SL}(2,A))\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{SL}(2,A))_{R_{S}}$
is the refined weight space decomposition of $\operatorname{Lie}(\operatorname{SL}(2,A))$.
3. The algebra of distributions of $\operatorname{GL}(2,A)$ and $\operatorname{SL}(2,A)$
In this section we include some definitions and results on the algebra of distributions $\operatorname{Dist}_{n}(G,e)$ of a group scheme $G$ over a commutative ring $k$. We also calculate the module of first order distributions and Lie algebras of $\operatorname{SL}(2,A)$ and $\operatorname{GL}(2,A)$. Let $(k[G],\Delta,S,\epsilon)$ be the Hopf algebra of the group scheme $G$ and let $I:=ker(\epsilon)$ be the kernel of the counit - the augmentation ideal.
@@ 3. The algebra of distributions of $\operatorname{GL}(2,A)$ and $\operatorname{SL}(2,A)$
Let $k[G]:=A[x_{ij}][1/D]$ be the localization of $A[x_{ij}]$ at the determinant $D$ and let $\operatorname{GL}(2,A):=\operatorname{Spec}(k[G])$. It follows for any commutative $A$-algebra $B$ the group $h_{k[G]}(B)$ is the group of $2\times 2$-matrices with coefficients in $B$ and determinant a unit in $B$. There is a canonical surjective map $k[G]\rightarrow k[S]$ with kernel $(D-1)$ hence the group scheme $\operatorname{SL}(2,A):=\operatorname{Spec}(k[S])$ is a closed subgroup scheme of $\operatorname{GL}(2,A)$. We wil calculate the module of distributions $\operatorname{Dist}_{1}(\operatorname{GL}(2,A),e)$ and $\operatorname{Dist}_{1}(\operatorname{SL}(2,A),e)$ and the Lie algebras $\operatorname{Lie}(\operatorname{GL}(2,A))$ and $\operatorname{Lie}(\operatorname{SL}(2,A))$.
@@
$\operatorname{Dist}_{1}(\operatorname{GL}(2,A),e)\cong A\{z_{0},z_{1},z_{2},z_{3},z_{4}\}.$
@@
and we define $[z_{i},z_{j}]:=\psi(z_{i},z_{j})\circ\rho^{\prime}\in\operatorname{Dist}_{1}(\operatorname{GL}(2,A),e)$. Let us calculate $[z_{2},z_{3}]$ as an example:
@@
Let $\mathfrak{gl}(2,A)$ denote the $A$-Lie algebra of $2\times 2$-matrices with coefficients in $A$. It follows $\mathfrak{gl}(2,A)\cong A\{t_{1},t_{2},\tilde{x},\tilde{y}\}$ is an isomorphism of $A$-Lie algebras. There is an isomorphism $\mathfrak{sl}(2,A)\cong A\{\tilde{x},\tilde{H},\tilde{y}\}$ with $\tilde{H}:=t_{1}-t_{2}$.
There is a surjection $R^{\prime}\rightarrow R$ with kernel $(D-1)$ giving a closed embedding of group schemes $\operatorname{SL}(2,A)\subseteq\operatorname{GL}(2,A)$ We get a canonical surjective map $p:R^{\prime}/(I^{\prime})^{2}\rightarrow R/I^{2}$ and an induced embedding of modules of distributions
$g:\operatorname{Dist}_{1}(\operatorname{SL}(2,A),e)\subseteq\operatorname{Dist}_{1}(\operatorname{GL}(2,A),e).$
The $A$-module $R/I^{2}$ has basis $R/I^{2}\cong A\{1,x_{12},x_{11}-1,x_{21}\}\cong A\{v_{0},e_{x},e_{H},e_{y}\}$ with $v_{0}:=1,e_{x}:=x_{12},e_{H}:=x_{11}-1$ and $e_{y}:=x_{21}$. The dual $\operatorname{Dist}_{1}(\operatorname{SL}(2,A),e)$ has basis
$\operatorname{Dist}_{1}(\operatorname{SL}(2,A),e)\cong A\{w_{0},x,H,y\}$
@@
The following holds: $[z_{0},z_{i}]=0$ for $i=1,2,3,4$. Moreover $[z_{1},z_{2}]=z_{2},[z_{1},z_{3}]=-z_{3},[z_{1},z_{4}]=0,[z_{2},z_{3}]=z_{1}-z_{4},[z_{2},z_{4}]=z_{2}$ and $[z_{3},z_{4}]=-z_{3}$. Hence there is an isomorphism $f:\operatorname{Dist}_{1}(\operatorname{GL}(2,A),e)\cong A\{z_{0}\}\oplus\mathfrak{gl}(2,A)$ defined by $f(z_{1}):=t_{1},f(z_{2}):=\tilde{x},f(z_{3}):=\tilde{y}$ and $f(z_{4}):=t_{2}$. The embedding $g$ is defined as follows: $g(w_{0}):=z_{0},g(x):=z_{2}=\tilde{x},g(H):=z_{1}-z_{4}=\tilde{H}$ and $g(y):=z_{3}=\tilde{y}$. Hence it identifies $\operatorname{Dist}_{1}(\operatorname{SL}(2,A),e)$ with the sub Lie algebra $A\{z_{0}\}\oplus\mathfrak{sl}(2,A)$ of $\operatorname{Dist}_{1}(\operatorname{GL}(2,A),e)$ under the isomorphism $f$.
@@
Hence the $A$-Lie algebras $\operatorname{Lie}(\operatorname{SL}(2,A))$ and $\operatorname{Lie}(\operatorname{GL}(2,A))$ are ”what we expect them to be”.
4. On the schematic normalizer of the diagonal torus in $\operatorname{SL}(2,A)$
In this section we study the group scheme $G:=\operatorname{SL}(2,A)$ for any commutative ring $A$ and the schematic normalizer $N\subseteq\operatorname{SL}(2,A)$ of the diagonal torus. The diagonal torus $T\subseteq\operatorname{SL}(2,A)$ is contained in the schematic normalizer $N$ and the schematic normalizer $N\subseteq N_{G}(T)$ is a closed sub group scheme of the functor normalizer $N_{G}(T)$ as defined in the book of Jantzen. There is an equality $N=N_{G}(T)$ if $A$ contains a field with more than $3$ elements. There are some inconsistencies in [8] in ths usage of $N_{G}(T)$ and its relation with the Weyl group, and for this reason we will use the schematic normalizer $N$ in this paper. In [8] it is claimed the normalizer $N_{G}(T)$ is the finite group scheme associated to the Weyl group and this is wrong in the case of $\operatorname{SL}(2,\mathbf{Z})$. We give an explicit construction of the quotient $N/T$ and its representing $\mathbf{Z}$-Hopf algebra $k[N/T]$, and use this construction to prove this.
@@ 4. On the schematic normalizer of the diagonal torus in $\operatorname{SL}(2,A)$
by $h_{R}(B):=\operatorname{Hom}_{A-alg}(R,B)$. It follows $h_{R}(B)$ is the group of $2\times 2$ matrices with coefficients in $B$ and determinant $1$. It follows $G:=\operatorname{SL}(2,A):=\operatorname{Spec}(R)$ is an affine group scheme in the sense of the book of Demazure/Gabriel and Jantzen and the ring $R$ is a Hopf algebra.
@@
It follows $R/I\cong R_{1}\oplus R_{2}$ with $R_{1}:=A[t,\frac{1}{t}]$ and $R_{2}:=A[u,v]/(uv+1)$. The schemes $T,N$ are closed subgroup schemes of $\operatorname{SL}(2,A)$. The group scheme $T$ is a closed and normal subgroup scheme of $N$.
@@
and it follows $T,N\subseteq\operatorname{SL}(2,A)$ are closed subgroup schemes. There is an equality $I_{1}+I_{2}=(1)$, hence the two ideals $I_{1},I_{2}$ are coprime. It follows $I:=I_{1}I_{2}=I_{1}\cap I_{2}$ annd by the Chinese remaider Lemma there is an isomorphism of rings
@@
The scheme $T$ is the diagonal torus in $\operatorname{SL}(2,A)$ and the scheme $N$ is the schematic normalizer of $T$.
There is the functor normalizer $N_{G}(T)$ as defined in the book of Jantzen and there is in general a strict inclusion $N\subsetneq N_{G}(T)$. There is an equality $N=N_{G}(T)$ when $A$ contains a field with more than 3 elements. If $A:=\mathbf{Z}$ is the ring of integers there is a strict inclusion $N\subsetneq N_{G}(T)$ of group schemes.
@@
Note: If $2\in A$ is invertible there is an isomorphism of Hopf algebras $k[N/T]\cong A[S(2)]$ where $A[S(2)]$ is the Hopf algebra on the symmetric group $S(2)$ on two elements. In this case it follows the group scheme $N/T$ is the finite group scheme (over $A$) associated to $S(2)$. The exact sequence
@@
and one checks $N_{\mathbf{C}}/T_{\mathbf{C}}\cong S(2)$ is the symmetric group on two letters. One may calculate the cohomology class $c(N_{\mathbf{C}})\in\operatorname{H}^{2}(S(2),T)$ associated to the above exact sequence and one will find the class $c(N_{\mathbf{C}})$ is non trivial. Hence the sequence does not split.
If $A:=\mathbf{Z}$ is the ring of integers there is no isomorphism $k[N/T]\cong\mathbf{Z}[S(2)]$. Hence $N/T$ is not the finite group scheme (over $\mathbf{Z}$) of the symmetric group $S(2)$ on two elements. The symmetric group $S(2)$ is the Weyl group of $\operatorname{SL}(2,\mathbf{Z})$.
Note: The following is claimed in the book of [8]: The functor normalizer $N_{G}(T)\subseteq\operatorname{SL}(2,\mathbf{Z})$ is a closed subgroup scheme and the quotient $N_{G}(T)/T$ equals the finite group scheme (over $\mathbf{Z}$) of the symmetric group $S(2)$. This cannot be true for the following reason: There is a closed immersion $N\subseteq N_{G}(T)$ and this closed immersion gives rise to a closed immersion $N/T\subseteq N_{G}(T)/T$. Such a closed immersion would give rise to a surjection
$\rho:\mathbf{Z}[S(2)]\rightarrow k[N/T]\cong\mathbf{Z}\oplus\mathbf{Z},$
but you may check there cannot exist a surjection $\rho:\mathbf{Z}[S(2)]\rightarrow\mathbf{Z}\oplus\mathbf{Z}$. Hence it appears there is a minor mistake in Jantzens book. For this reason we willl use the schematic normalizer in this paper.
If $V:=A^{n}$ is a free rank $n$ $A$-module and if $G:=\operatorname{SL}(V)$ is the ”special linear group scheme” on $V$, we may define the ”diagonal torus” $T\subseteq\operatorname{SL}(V)$. The diagonal torus $T$ is a closed subgroup scheme of $\operatorname{SL}(V)$ and we may define its ”functor normalizer” $N_{G}(T)$ in the sense of [8]. We may define the schematic normalizer $N$ in this more general situation, and $N$ is a closed subgroup scheme of $N_{G}(T)$ and these two group schemes agree when $A$ contains a field with more than $3$ elements.
There is always an exact sequence of group schemes
$e\rightarrow T\rightarrow N\rightarrow N/T\rightarrow e$
and $N/T$ is an affine group scheme of finite type over $A$. There is in general a closed embedding $N/T\subseteq N_{G}(T)/T$.
5. On the schematic normalizer in $\operatorname{GL}(2,A)$
In this section we study the group scheme $G:=\operatorname{GL}(2,A)$ where $A$ is an arbitrary commutative unital ring, the diagonal torus $T\subseteq G$ and the schematic normalizer $N\subseteq G$ of $T$. We prove there is an exact sequence of affine group schemes

### v4

@@ 2. The adjoint representation of $\operatorname{GL}_{2,A}$ and $\operatorname{SL}_{2,A}$
The adjoint representation for $\operatorname{Lie}(\operatorname{SL}_{2,A})$.
Let $\mathfrak{sl}_{2,A}:=A\{e_{12},e_{11}-e_{22},e_{21}\}\subseteq\mathfrak{gl}_{2,A}$, and let $x:=e_{12},H:=e_{11}-e_{22},y:=e_{21}$ and let $B(s):=\{x,H,y\}$. Define for any commutative $A$-algebra $C$ the following: Let $B(s)_{C}:=\{x\otimes 1_{C},H\otimes 1_{C},y\otimes 1_{C}\}$ be a basis for $\mathfrak{sl}_{2,A}\otimes_{A}C$. Here $1_{C}\in C$ is the multiplicative unit. You may check the following:
@@
Hence $\mathfrak{sl}_{2,A}\subseteq\mathfrak{gl}_{2,A}$ is a sub $k[G]$-comodule. Let $S:=\operatorname{SL}_{2,A}:=V(D-1)\subseteq\operatorname{GL}_{2,A}$. we get the induced $k[S]$-comodule
$\operatorname{Ad}(S):\mathfrak{sl}_{2,A}\rightarrow\mathfrak{sl}_{2,A}\otimes_{A}k[S]$
@@
There is an isomorphism $\operatorname{Lie}(\operatorname{SL}_{2,A})\cong\mathfrak{sl}_{2,A}$. Let $\operatorname{Lie}(T):=A\{H\}$ and let $\operatorname{Lie}(\operatorname{SL}_{2,A})_{R_{S}}:=A\{x,y\}$. It follow there are comodule structures
@@
$\Delta_{2}:\operatorname{Lie}(\operatorname{SL}_{2,A})_{R_{S}}\rightarrow\operatorname{Lie}(\operatorname{SL}_{2,A})_{R_{S}}\otimes k[N]$
@@
$\operatorname{Lie}(\operatorname{SL}_{2,A})\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{SL}_{2,A})_{R_{S}}.$
@@
The set $R_{S}:=\{-2,2\}$ is the roots of $\operatorname{SL}_{2,A}$ and the direct sum decomposition
$\operatorname{Lie}(\operatorname{SL}_{2,A})\cong\operatorname{Lie}(T)\oplus\operatorname{Lie}(\operatorname{SL}_{2,A})_{R_{S}}$
is the refined weight space decomposition of $\operatorname{Lie}(\operatorname{SL}_{2,A})$.
3. The algebra of distributions of $\operatorname{GL}_{2,A}$ and $\operatorname{SL}_{2,A}$
In this section we include some definitions and results on the algebra of distributions $\operatorname{Dist}_{n}(G,e)$ of a group scheme $G$ over a commutative ring $k$. We also calculate the module of first order distributions and Lie algebras of $\operatorname{SL}_{2,A}$ and $\operatorname{GL}_{2,A}$. Let $(k[G],\Delta,S,\epsilon)$ be the Hopf algebra of the group scheme $G$ and let $I:=ker(\epsilon)$ be the kernel of the counit - the augmentation ideal.
@@ 3. The algebra of distributions of $\operatorname{GL}_{2,A}$ and $\operatorname{SL}_{2,A}$
Let $k[G]:=A[x_{ij}][1/D]$ be the localization of $A[x_{ij}]$ at the determinant $D$ and let $\operatorname{GL}_{2,A}:=\operatorname{Spec}(k[G])$. It follows for any commutative $A$-algebra $B$ the group $h_{k[G]}(B)$ is the group of $2\times 2$-matrices with coefficients in $B$ and determinant a unit in $B$. There is a canonical surjective map $k[G]\rightarrow k[S]$ with kernel $(D-1)$ hence the group scheme $\operatorname{SL}_{2,A}:=\operatorname{Spec}(k[S])$ is a closed subgroup scheme of $\operatorname{GL}_{2,A}$. We wil calculate the module of distributions $\operatorname{Dist}_{1}(\operatorname{GL}_{2,A},e)$ and $\operatorname{Dist}_{1}(\operatorname{SL}_{2,A},e)$ and the Lie algebras $\operatorname{Lie}(\operatorname{GL}_{2,A})$ and $\operatorname{Lie}(\operatorname{SL}_{2,A})$.
@@
$\operatorname{Dist}_{1}(\operatorname{GL}_{2,A},e)\cong A\{z_{0},z_{1},z_{2},z_{3},z_{4}\}.$
@@
and we define $[z_{i},z_{j}]:=\psi(z_{i},z_{j})\circ\rho^{\prime}\in\operatorname{Dist}_{1}(\operatorname{GL}_{2,A},e)$. Let us calculate $[z_{2},z_{3}]$ as an example:
@@
Let $\mathfrak{gl}_{2,A}$ denote the $A$-Lie algebra of $2\times 2$-matrices with coefficients in $A$. It follows $\mathfrak{gl}_{2,A}\cong A\{t_{1},t_{2},\tilde{x},\tilde{y}\}$ is an isomorphism of $A$-Lie algebras. There is an isomorphism $\mathfrak{sl}_{2,A}\cong A\{\tilde{x},\tilde{H},\tilde{y}\}$ with $\tilde{H}:=t_{1}-t_{2}$.
There is a surjection $R^{\prime}\rightarrow R$ with kernel $(D-1)$ giving a closed embedding of group schemes $\operatorname{SL}_{2,A}\subseteq\operatorname{GL}_{2,A}$ We get a canonical surjective map $p:R^{\prime}/(I^{\prime})^{2}\rightarrow R/I^{2}$ and an induced embedding of modules of distributions
$g:\operatorname{Dist}_{1}(\operatorname{SL}_{2,A},e)\subseteq\operatorname{Dist}_{1}(\operatorname{GL}_{2,A},e).$
The $A$-module $R/I^{2}$ has basis $R/I^{2}\cong A\{1,x_{12},x_{11}-1,x_{21}\}\cong A\{v_{0},e_{x},e_{H},e_{y}\}$ with $v_{0}:=1,e_{x}:=x_{12},e_{H}:=x_{11}-1$ and $e_{y}:=x_{21}$. The dual $\operatorname{Dist}_{1}(\operatorname{SL}_{2,A},e)$ has basis
$\operatorname{Dist}_{1}(\operatorname{SL}_{2,A},e)\cong A\{w_{0},x,H,y\}$
@@
The following holds: $[z_{0},z_{i}]=0$ for $i=1,2,3,4$. Moreover $[z_{1},z_{2}]=z_{2},[z_{1},z_{3}]=-z_{3},[z_{1},z_{4}]=0,[z_{2},z_{3}]=z_{1}-z_{4},[z_{2},z_{4}]=z_{2}$ and $[z_{3},z_{4}]=-z_{3}$. Hence there is an isomorphism $f:\operatorname{Dist}_{1}(\operatorname{GL}_{2,A},e)\cong A\{z_{0}\}\oplus\mathfrak{gl}_{2,A}$ defined by $f(z_{1}):=t_{1},f(z_{2}):=\tilde{x},f(z_{3}):=\tilde{y}$ and $f(z_{4}):=t_{2}$. The embedding $g$ is defined as follows: $g(w_{0}):=z_{0},g(x):=z_{2}=\tilde{x},g(H):=z_{1}-z_{4}=\tilde{H}$ and $g(y):=z_{3}=\tilde{y}$. Hence it identifies $\operatorname{Dist}_{1}(\operatorname{SL}_{2,A},e)$ with the sub Lie algebra $A\{z_{0}\}\oplus\mathfrak{sl}_{2,A}$ of $\operatorname{Dist}_{1}(\operatorname{GL}_{2,A},e)$ under the isomorphism $f$.
@@
Hence the $A$-Lie algebras $\operatorname{Lie}(\operatorname{SL}_{2,A})$ and $\operatorname{Lie}(\operatorname{GL}_{2,A})$ are ”what we expect them to be”.
4. On the schematic normalizer of the diagonal torus in $\operatorname{SL}_{2,A}$
In this section we study the group scheme $G:=\operatorname{SL}_{2,A}$ for any commutative ring $A$ and the schematic normalizer $N\subseteq\operatorname{SL}_{2,A}$ of the diagonal torus. The diagonal torus $T\subseteq\operatorname{SL}_{2,A}$ is contained in the schematic normalizer $N$ and the schematic normalizer $N$ is isomorphic to the functor normalizer $N_{G}(T)$ as defined in the book of Jantzen [8].
@@ 4. On the schematic normalizer of the diagonal torus in $\operatorname{SL}_{2,A}$
by $h_{R}(B):=\operatorname{Hom}_{A-alg}(R,B)$. It follows $h_{R}(B)$ is the group of $2\times 2$ matrices with coefficients in $B$ and determinant $1$. It follows $G:=\operatorname{SL}_{2,A}:=\operatorname{Spec}(R)$ is an affine group scheme in the sense of the book of Demazure/Gabriel and Jantzen and the ring $R$ is a Hopf algebra.
@@
It follows $R/I\cong R_{1}\oplus R_{2}$ with $R_{1}:=A[t,\frac{1}{t}]$ and $R_{2}:=A[u,v]/(uv+1)$. The schemes $T,N$ are closed subgroup schemes of $\operatorname{SL}_{2,A}$. The group scheme $T$ is a closed and normal subgroup scheme of $N$.
@@
and it follows $T,N\subseteq\operatorname{SL}_{2,A}$ are closed subgroup schemes. There is an equality $I_{1}+I_{2}=(1)$, hence the two ideals $I_{1},I_{2}$ are coprime. It follows $I:=I_{1}I_{2}=I_{1}\cap I_{2}$ annd by the Chinese remaider Lemma there is an isomorphism of rings
@@
The scheme $T$ is the diagonal torus in $\operatorname{SL}_{2,A}$ and the scheme $N$ is the schematic normalizer of $T$.
Proposition 4.3.
There is an isomorphism of group schemes $N\cong N_{G}(T)$ where $N_{G}(T)$ is the functor normalizer as defined in [8].
Proof.
Let $T\subseteq S:=\operatorname{SL}_{2,A}$ be the diagonal torus. By definition it follows $T:=V(I_{1})$ is defined by the ideal $I_{1}$. There is an equality of ideals $I=(x_{11}x_{12},x_{21}x_{22})$. By definition it follows $h_{N_{G}(T)}(R)$ is the group of all elelments $g\in h_{S}(R)$ with the property that for any map $f:R\rightarrow R_{1}$ and any element $t\in h_{T}(R_{1})$, it follows $gtg^{-1}\in h_{T}(R_{1})$. If $g\in V(I_{1})$ it follows $g\in h_{N_{G}(T)}(R)$. Conversely if $g\in h_{N_{G}(T)}(R)$ the following holds for any element $t\in h_{T}(R_{1})$ with $R_{1}:=R[z,1/z]$: $gtg^{-1}\in h_{T}(R_{1})$. Let $g:=(a_{ij})$ be the 2 times 2 matrix with coefficients in $R$ and determinant $D:=a_{11}a_{22}-a_{12}a_{21}$ a unit in $R$ corresponding to $g$. Let $t\in h_{T}(R_{1})$ be the element
$t:=\begin{pmatrix}z&0\\ 0&z^{-1}\end{pmatrix}.$
It follows
$gtg^{-1}:=\begin{pmatrix}(a_{11}a_{22}z-a_{12}a_{21}/z)\frac{1}{D}&a_{11}a_{12}(1/z-z)\frac{1}{D}\\ a_{21}a_{22}(z-1/z)\frac{1}{D}&(a_{11}a_{22}/z-a_{12}a_{21}z)\frac{1}{D}\end{pmatrix}.$
If $gtg^{-1}\in h_{T}(R_{1})$ it follows $a_{11}a_{12}=a_{21}a_{22}=0$ and this implies $g\in V(I_{1})$. Hence $g\in h_{N}(A)$ and the Proposition follows. ∎
@@
The exact sequence
@@
and one checks $N_{\mathbf{C}}/T_{\mathbf{C}}\cong S(2)$ is the symmetric group on two letters. One may calculate the cohomology class $c(N_{\mathbf{C}})\in\operatorname{H}^{2}(S(2),T)$ associated to the above exact sequence and one will find the class $c(N_{\mathbf{C}})$ is non trivial. Hence the sequence does not split. The Weyl group $W$ of $\operatorname{SL}(2,\mathbf{C})$ is isomorphic to $S(2)$.
5. On the schematic normalizer in $\operatorname{GL}_{2,A}$
In this section we study the group scheme $G:=\operatorname{GL}_{2,A}$ where $A$ is an arbitrary commutative unital ring, the diagonal torus $T\subseteq G$ and the schematic normalizer $N\subseteq G$ of $T$. We prove there is a split exact sequence of affine group schemes